# NEAT: Null-Space Constrained Model Editing in A Transfromer

### v2

@@ 6 Conclusion
In this work, we introduced AlphaEdit, a novel model editing method to address a critical challenge in current approaches — the trade-off between knowledge update and preservation — with only a single line of code. Specifically, AlphaEdit minimizes disruption to the preserved knowledge by projecting parameter perturbations onto the null space of its key matrices. It then removes the output error related to it from the current objective, allowing the model to focus solely on knowledge update without trade-off. Extensive experiments on multiple base LLMs, including LLaMA3, GPT-2 XL, and GPT-J, demonstrate that AlphaEdit significantly enhances the performance of existing model editing methods, delivering an average improvement of $36.4\%$ in editing capabilities.
@@ A.2 Metrics
Now we introduce the evaluation metrics used for the above two datasets, respectively.
@@ A.3 Implementation Details
All experiments are conducted on a single A40 (48GB) GPU. The LLMs are loaded using HuggingFace Transformers (Wolf et al., 2019). For GPT-2 XL, the average editing time per sample is 2.74 seconds, while for GPT-J, it is 6.18 seconds.
@@ A.4 Baselines
Appendix B Model Editing & Related Proofs
Here, we provide the specific implementation details of the current model editing algorithm (Meng et al., 2023) along with the proof process related to it and the concept of the null space (Wang et al., 2021).
@@ B.1 Model Editing
Model editing aims to refine a pre-trained model through one or multiple edits, while each edit replaces $(s,r,o)$ with the new knowledge $(s,r,o^{\ast})$. Then, model is expected to recall the updated object $o^{\ast}$ given a natural language prompt $p(s,r)$, such as “The President of the United States is.”
To achieve this, locating-and-editing methods are proposed for effectively model editing (Meng et al., 2022; 2023). These methods typically adhere to the following steps.
@@ B.3 Proof for Equation $\Delta PK_{0}(K_{0})^{T}=0$
Based on the Eqn. equation 28 and 29, we can derive that:
@@ B.4 Invertibility of $(K_{p}K_{p}^{T}P+K_{1}K_{1}^{T}P+\alpha I)$
To prove the invertibility of the matrix $({\bm{K}}_{p}{\bm{K}}_{p}^{T}P+{\bm{K}}_{1}{\bm{K}}_{1}^{T}{\bm{P}}+\alpha{\bm{I}})$, note that ${\bm{K}}_{p}{\bm{K}}_{p}^{T}$ and ${\bm{K}}_{1}{\bm{K}}_{1}^{T}$ are symmetric and positive semidefinite matrices, and ${\bm{P}}$ is a projection matrix which is also symmetric and positive semidefinite. Since ${\bm{P}}$ projects onto a subspace, the matrices ${\bm{K}}_{p}{\bm{K}}_{p}^{T}{\bm{P}}$ and ${\bm{K}}_{1}{\bm{K}}_{1}^{T}{\bm{P}}$ are positive semidefinite.
@@ C.3 Quantification of Distribution Shift
According to the results in Figure 9, we observe that across all metrics, baseline methods, and base LLMs, the methods optimized with AlphaEdit consistently exhibit minimal distributional shifts. This further supports the qualitative analysis presented in the main text, demonstrating that AlphaEdit effectively prevents post-edited LLMs from overfitting hidden representations to the updated knowledge, thereby preserving the original knowledge.
Figure 9: The quantitative results of hidden representation shifts before and after editing Best viewed in color.
@@ C.4 Editing Facts involving Various Semantics
To gain a deeper understanding of the performance when editing facts involving different semantics, following MEMIT (Meng et al., 2023), we selected several semantics from the Counterfact dataset, each containing at least 300 examples, and evaluated the performance of each baseline methods on these examples (which were evenly distributed across the sequential editing batches). Some of the results are shown in Figure 7 in the main text, with more comprehensive results displayed in Figure 10. In these figures, the horizontal axis represents Accuracy, defined as the average of Efficacy and Generalization across the 300 examples. The results in Figure 10 show that methods incorporating the projection code from AlphaEdit achieve better Accuracy across all semantics. It also indicates that some semantic categories are more challenging to edit than others. However, even in these more difficult cases, baseline methods enhanced with AlphaEdit’s projection code still achieve over 90% accuracy.
@@
Appendix D Limitations & Future Discussion
While we acknowledge the effectiveness of our method, we also recognize its limitations. Concretely, our method has been tested exclusively within the locating-and-editing paradigm, but we have not extended it to other frameworks, such as parameter-preserving methods. Additionally, our experimental results indicate that in cases where the number of edits is very small, baselines perform sufficiently well (i.e., causing minimal degradation to the model). As a result, AlphaEdit’s advantages are less pronounced in these scenarios. Moving forward, future research could focus on extending AlphaEdit to a broader range of editing paradigms and exploring its potential for enhancing single-edit performance. These directions offer promising avenues for improving both the generality and impact of our approach. Furthermore, we plan to extract the null space for “difficult samples” (Hu et al., 2024) utilizing techniques such as explainability (Fang et al., 2023) and anomaly detection Gao et al. (2024) to optimize the knowledge updating.
Appendix E Visualizing the ZSRE and Counterfact Datasets Through Examples

### v3

@@ 7 Conclusion
In this work, we introduced AlphaEdit, a novel model editing method to address a critical challenge in current approaches — the trade-off between knowledge update and preservation — with only a single line of code. Specifically, AlphaEdit minimizes disruption to the preserved knowledge by projecting parameter perturbations onto the null space of its key matrices, allowing the model to focus solely on knowledge update. Extensive experiments on multiple base LLMs, including LLaMA3, GPT-2 XL, and GPT-J, demonstrate that AlphaEdit significantly enhances the performance of existing model editing methods, delivering an average improvement of $36.7\%$ in editing capabilities.
@@
Acknowledgement
This research is supported by the National Science and Technology Major Project (2023ZD0121102), the National Natural Science Foundation of China (92270114), and the National Natural Science Foundation of China (U24B20180). We also thank the EasyEdit platform for their support (https://github.com/zjunlp/EasyEdit).
@@ A.1 Datasets
Here, we provide a detailed introduction to the datasets used in this paper:
@@
  • KnowEdit (Zhang et al., 2024d) introduces a comprehensive benchmark aimed at systematically evaluating knowledge editing methods, categorizing them into approaches that rely on external knowledge, intrinsic knowledge updates, or merging new knowledge into the model. The benchmark not only measures the impact of editing on specific domains but also emphasizes preserving the model’s overall performance across tasks, offering a unified framework for evaluating editing efficiency and impact. In our paper, we employ the wiki_recent and wikibio within KnowEdit to conduct our experiments.
  • LEME (Long-form Evaluation of Model Editing) (Rosati et al., 2024) extends the evaluation paradigm by focusing on long-form generative outputs, revealing unique challenges such as factual drift, internal consistency, and lexical cohesion. This protocol highlights that short-form metrics fail to correlate with long-form generative outcomes, shedding light on previously unexplored dimensions of editing.
  • MQuAKE (Zhong et al., 2023) addresses a critical gap in current evaluations by introducing multi-hop reasoning questions to test the ripple effects of factual updates. Unlike single-fact recall benchmarks, MQUAKE measures the consistency of entailed beliefs after editing, uncovering limitations in existing methods when handling complex relational dependencies.
@@ A.2 Metrics
Now we introduce the evaluation metrics used for ZsRE and Counterfact datasets, respectively.
@@ A.3 Implementation Details
All experiments are conducted on a single A40 (48GB) GPU. The LLMs are loaded using HuggingFace Transformers (Wolf et al., 2019).
@@ A.4 Baselines
  • InstructEdit enables the learning of a well-formed Editor by designing corresponding instructions for training on different tasks. InstructEdit applies meta-learning editing methods based on MEND to train the editor with a variety of meticulously curated instructions, and through this approach, InstructEdit can endow the Editor with the capacity for multi-task editing, thus saving a significant amount of human and computational resources.
@@
  • SERAC (Mitchell et al., 2022b) introduces Semi-Parametric Editing with a Retrieval-Augmented Counterfactual Model, addressing limitations of traditional editors in defining edit scope and handling sequential updates. It stores edits in explicit memory and reasons over them to adjust model behavior as needed. SERAC outperforms existing methods on three challenging tasks—question answering, fact-checking, and dialogue generation.
  • MELO (Yu et al., 2024) proposes Neuron-Indexed Dynamic LoRA, a plug-in method that dynamically activates LoRA blocks to edit model behavior efficiently. Adaptable across multiple LLM backbones, it achieves state-of-the-art performance on tasks like document classification, question answering, and hallucination correction, with minimal computational cost and trainable parameters.
  • GRACE (Hartvigsen et al., 2023) introduces a lifelong model editing framework that uses a local codebook in the latent space to handle thousands of sequential edits without degrading model performance. It makes targeted fixes while preserving generalization, demonstrating superior results on T5, BERT, and GPT for retaining and generalizing edits.
Appendix B Implementation Details of Current Model Editing & Related Proofs
Here, we provide the implementation details of the current model editing methods along with the proof process related to it and the concept of the null space (Wang et al., 2021).
@@ B.1 Model Editing
Model editing aims to refine a pre-trained model through one or multiple edits, while each edit replaces $(s,r,o)$ with the new knowledge $(s,r,o^{\ast})$ (Yang et al., 2024b; Li et al., 2025). Then, model is expected to recall the updated object $o^{\ast}$ given a natural language prompt $p(s,r)$, such as “The President of the United States is” (Zhang et al., 2024c).
To achieve this, locating-and-editing methods are proposed for effectively model editing (Wang et al., 2024; Yang et al., 2024a). These methods typically adhere to the following steps (Jiang et al., 2024):
@@ B.3 Proof for Equation $\Delta PK_{0}(K_{0})^{T}=0$
Based on the Eqn. 28 and 29, we can derive that:
@@ B.4 Invertibility of $(K_{p}K_{p}^{T}P+K_{1}K_{1}^{T}P+\alpha I)$
To prove the invertibility of the matrix $({\bm{K}}_{p}{\bm{K}}_{p}^{T}{\bm{P}}+{\bm{K}}_{1}{\bm{K}}_{1}^{T}{\bm{P}}+\alpha{\bm{I}})$, note that ${\bm{K}}_{p}{\bm{K}}_{p}^{T}$ and ${\bm{K}}_{1}{\bm{K}}_{1}^{T}$ are symmetric and positive semidefinite matrices, and ${\bm{P}}$ is a projection matrix which is also symmetric and positive semidefinite. Since ${\bm{P}}$ projects onto a subspace, the matrices ${\bm{K}}_{p}{\bm{K}}_{p}^{T}{\bm{P}}$ and ${\bm{K}}_{1}{\bm{K}}_{1}^{T}{\bm{P}}$ are positive semidefinite.
@@ C.3 Quantification of Distribution Shift
According to the results in Figure 9, we observe that across all metrics and base LLMs, the methods optimized with AlphaEdit consistently exhibit minimal distributional shifts. This further supports the qualitative analysis presented in the main text, demonstrating that AlphaEdit effectively prevents post-edited LLMs from overfitting hidden representations to the updated knowledge, thereby preserving the original knowledge.
Figure 9: The quantitative results of hidden representation shifts before and after editing. Best viewed in color.
@@ C.4 Editing Facts involving Various Semantics
To gain a deeper understanding of the performance when editing facts involving different semantics, following MEMIT (Meng et al., 2023), we selected several semantics from the Counterfact dataset, each containing at least 300 examples, and evaluated the performance of each baseline methods on these examples (which were evenly distributed across the sequential editing batches). Some of the results are shown in Figure 7 in the main text, with more comprehensive results displayed in Figure 10. In these figures, the horizontal axis represents Accuracy, defined as the average of Efficacy and Generalization across the 300 examples. For instance, the bar labeled "language," with a height of 98, indicates that out of 1,000 knowledge instances related to language (e.g., "The primary language in the United States is English," or "The official language of France is French"), AlphaEdit successfully edited 98% of them. This metric provides a fine-grained assessment of AlphaEdit’s effectiveness across various knowledge domains. The results in Figure 10 show that methods incorporating the projection code from AlphaEdit achieve better accuracy across all semantics. It also indicates that some semantics are more challenging to edit than others. However, even in these more difficult cases, baseline methods enhanced with AlphaEdit’s projection code still achieve over 90% accuracy.
@@
C.5 Comparison between AlphaEdit and Memory-based Editing Methods
Although our AlphaEdit mainly focuses on the parameter-modifying editing method, we also hope to explore the advantages of our method compared with some mainstream memory-based editing methods. Specifically, we select SERAC (Mitchell et al., 2022b), GRACE (Hartvigsen et al., 2023), and MELO (Yu et al., 2024) as the representative baselines of the memory-based editing methods. The results are shown in Table 2. According to Table 2 we can summarize that:
  • Across all models and tasks, AlphaEdit consistently achieves the highest scores in efficacy (Eff.) and generalization (Gen.). This indicates that AlphaEdit is highly effective at correctly applying the desired edits while maintaining robust generalization to the other knowledge. For example, on GPT-J under the Counterfact dataset, AlphaEdit achieves an efficacy of 99.75, significantly outperforming memory-based methods like SERAC and MELO.
  • While AlphaEdit generally achieves competitive scores in specificity (Spe.) and fluency (Flu.), it does not always surpass the memory-based methods. However, we believe this trade-off is reasonable and acceptable because memory-based methods inherently rely on consuming storage space to better preserve existing knowledge.
Method Model Counterfact ZsRE
SERAC 71.21$\pm$0.56 61.05$\pm$0.39 66.90$\pm$0.21 615.72$\pm$0.34 20.77$\pm$0.13 67.75$\pm$0.24 33.96$\pm$0.35 22.17$\pm$0.15
GRACE 96.72$\pm$0.13 50.14$\pm$0.01 72.23$\pm$0.21 620.43$\pm$0.63 23.79$\pm$0.23 93.58$\pm$0.31 1.03$\pm$0.06 31.86$\pm$0.12
MELO 65.29$\pm$0.13 58.58$\pm$0.32 63.36$\pm$0.37 608.98$\pm$0.82 22.18$\pm$0.04 25.18$\pm$0.14 24.14$\pm$0.23 30.36$\pm$0.75
SERAC 82.28$\pm$0.26 58.31$\pm$0.34 68.98$\pm$0.32 615.92$\pm$0.72 28.65$\pm$0.17 92.37$\pm$0.29 38.21$\pm$0.32 25.17$\pm$0.25
GRACE 96.50$\pm$0.24 50.10$\pm$0.01 74.42$\pm$0.43 620.56$\pm$0.79 31.55$\pm$0.25 96.54$\pm$0.21 0.40$\pm$0.02 24.78$\pm$0.21
MELO 78.29$\pm$0.24 60.52$\pm$0.32 66.80$\pm$0.52 610.82$\pm$0.44 24.31$\pm$0.24 82.24$\pm$0.07 32.88$\pm$0.03 26.65$\pm$0.06
SERAC 72.25$\pm$0.15 58.18$\pm$0.23 64.06$\pm$0.37 595.35$\pm$0.35 27.35$\pm$0.12 92.17$\pm$0.67 36.57$\pm$0.72 20.67$\pm$0.22
GRACE 98.88$\pm$0.28 50.05$\pm$0.01 72.07$\pm$0.24 620.21$\pm$0.49 28.53$\pm$0.15 94.33$\pm$0.37 1.59$\pm$0.03 27.63$\pm$0.43
MELO 72.62$\pm$0.58 53.63$\pm$0.42 63.25$\pm$0.62 588.57$\pm$0.65 23.58$\pm$0.33 93.54$\pm$0.03 45.25$\pm$0.02 23.45$\pm$0.24
C.6 Evaluation on Additional Base LLMs: Gemma and phi-1.5
To enhance evaluation diversity, we extended experiments to two additional base LLMs, Gemma (Mesnard et al., 2024) and phi-1.5 (Li et al., 2023). We summarize the results in Table 3. These results demonstrate that AlphaEdit consistently outperforms MEMIT and RECT across key metrics on both the Counterfact and ZsRE datasets. Notably, on Gemma, AlphaEdit achieves the highest fluency (398.96) and consistency (32.91), reflecting its ability to maintain coherence and accuracy. Similarly, on phi-1.5, AlphaEdit excels in efficacy (70.79) and fluency (399.47), showcasing its adaptability to smaller, efficient models. These findings confirm AlphaEdit’s robustness across diverse LLM architectures and its capability to deliver high-quality edits while preserving model integrity.
Method Model Counterfact ZsRE
MEMIT Gemma 64.68$\pm$0.21 60.36$\pm$0.30 46.73$\pm$0.62 373.94$\pm$1.12 22.14$\pm$0.31 64.38$\pm$0.26 66.12$\pm$0.46 24.52$\pm$0.38
RECT 65.17$\pm$0.19 57.48$\pm$0.64 52.54$\pm$0.54 388.77$\pm$0.44 23.37$\pm$0.39 67.18$\pm$0.50 64.12$\pm$0.47 20.02$\pm$0.47
AlphaEdit 75.21$\pm$0.09 67.83$\pm$0.63 52.63$\pm$0.49 398.96$\pm$0.39 32.91$\pm$0.35 75.91$\pm$0.42 68.12$\pm$0.67 23.50$\pm$0.56
MEMIT phi-1.5 55.71$\pm$0.63 56.58$\pm$0.78 35.41$\pm$0.99 368.57$\pm$1.26 19.79$\pm$0.31 54.41$\pm$0.78 52.47$\pm$0.89 20.98$\pm$0.58
RECT 58.19$\pm$0.73 58.92$\pm$0.76 38.46$\pm$0.92 362.94$\pm$1.44 19.88$\pm$0.37 55.15$\pm$0.72 53.64$\pm$0.83 18.58$\pm$0.65
AlphaEdit 70.79$\pm$0.56 65.12$\pm$0.88 48.96$\pm$0.96 399.47$\pm$0.67 25.98$\pm$0.48 70.02$\pm$0.85 63.19$\pm$0.72 20.69$\pm$0.73
C.7 Evaluation on Expanding Benchmark: KnowEdit, LEME and MQUAKE
In this part, we selected two datasets from the KnowEdit (Zhang et al., 2024d) benchmark, namely wiki_recent and wikibio, for testing. During our experiments, we noticed that some samples within these datasets exhibit abnormally high norms in the hidden states when processed by the LLaMA3-8B-instruct model. These elevated norms are often accompanied by disproportionately large target value norms, which, if forcibly edited, could compromise the model’s stability. To mitigate this issue, we implemented an automatic filtering mechanism to exclude samples with excessively high hidden state norms during the continuous editing process. Experimental results are presented in Table 4.
Additionally, we extended our evaluations to two critical datasets, LEME (Rosati et al., 2024) and MQUAKE (Zhong et al., 2023), to test AlphaEdit’s performance in more challenging scenarios. LEME focuses on long-form generative tasks, emphasizing consistency, factual correctness, and lexical cohesion. MQUAKE, on the other hand, evaluates the ripple effects of factual updates using multi-hop reasoning questions. Results for these two datasets are summarized in Table 5.
According to Table 4 and 5 we can find that:
  • AlphaEdit demonstrates a remarkable ability to achieve high editing success rates across all evaluated datasets. For instance, on the wiki_recent dataset, AlphaEdit achieves an impressive 96.10% editing success, which is significantly higher than the second-best method, RECT (82.47%). A similar trend is observed on wikibio, where AlphaEdit reaches 95.34%, outperforming RECT by a substantial margin.
  • AlphaEdit demonstrates the best performance in both Multi-hop and Multi-hop (CoT) tasks, with scores of 9.14 and 9.75 respectively, significantly surpassing competing methods. This showcases its strong capability in handling complex reasoning tasks and ensuring logical consistency across interdependent facts.
  • On LEME, AlphaEdit excels in all three metrics, showcasing its ability to generate accurate long-form outputs. Its consistently high performance across GPT-J and GPT2-XL reinforces its reliability in executing precise edits while maintaining the structural coherence and integrity of the generated text.
Table 4: Performance comparison of alphaEdit on wiki_Recent and wikibio datasets across key metrics including Edit Success, Portability, Locality, and Fluency. Results are averaged with standard deviations. The best results are highlighted in bold.
Method wiki_recent wikibio
[…]
MEMIT 56.25$\pm$0.28 42.73$\pm$0.27 41.02$\pm$0.20 513.35$\pm$3.47 63.73$\pm$0.40 64.27$\pm$0.41 582.38$\pm$3.34
RECT 82.47$\pm$0.53 51.28$\pm$0.25 48.84$\pm$0.24 568.62$\pm$3.71 91.48$\pm$0.48 72.83$\pm$0.44 612.04$\pm$4.29
AlphaEdit 96.10$\pm$0.47 57.30$\pm$0.38 54.76$\pm$0.30 594.52$\pm$3.91 95.34$\pm$0.46 75.34$\pm$0.50 618.35$\pm$4.22
Table 5: Performance of AlphaEdit on MQuAKE and LEME datasets for multi-Hop reasoning and long-form editing. Metrics include Multi-hop Reasoning, Chain-of-Thought (CoT) Multi-hop Reasoning, Edit Consistency, Factual Consistency, and Internal Consistency. The best results are highlighted in bold.
Model Method MQuAKE LEME
[…]
MEMIT 3.35$\pm$0.07 6.13$\pm$0.12 2.11$\pm$0.18 2.02$\pm$0.17 3.84$\pm$0.29
GPT-J RECT 3.77$\pm$0.04 7.61$\pm$0.20 2.24$\pm$0.20 2.62$\pm$0.19 4.07$\pm$0.31
AlphaEdit 5.03$\pm$0.16 9.14$\pm$0.21 3.34$\pm$0.26 3.80$\pm$0.28 5.42$\pm$0.41
MEMIT 3.14$\pm$0.08 6.25$\pm$0.11 1.92$\pm$0.22 2.31$\pm$0.20 3.85$\pm$0.34
GPT2-XL RECT 3.72$\pm$0.06 7.48$\pm$0.24 2.12$\pm$0.26 2.60$\pm$0.21 4.13$\pm$0.29
AlphaEdit 5.00$\pm$0.23 9.25$\pm$0.27 3.28$\pm$0.36 3.07$\pm$0.33 5.76$\pm$0.49
C.8 Impact of Dataset Size on AlphaEdit’s Performance
To explore the relationship between dataset size for calculating ${\bm{K}}_{0}$ and AlphaEdit’s performance, we conduct additional experiments to evaluate the robustness of the model under reduced dataset conditions. Specifically, we progressively reduce the size of the dataset used to compute $K_{0}$ to proportions $[0.9,0.8,0.7,\dots,0.1]$ of its original size. The goal is to analyze the impact of dataset size on three key metrics: Efficacy, Generalization, and Specificity. All the results are summarized in Table 6. To further illustrate trend changes, we select the results for LLaMA3 and visualized them using line charts, as presented in Figure 11. According to Figure 11 we can find that:
  • As the dataset size decreased, both Efficacy and Generalization demonstrate notable stability. Even at only 10% of the original dataset size, the drop in these metrics is negligible (less than 5%), suggesting that AlphaEdit effectively generalizes to unseen data and remains efficient even with reduced data availability.
  • In contrast, the Specificity metric experience a significant decline as the dataset size is reduced. When the dataset size is limited to just 10% of its original volume, Specificity drop by 11.76%, indicating that the model’s ability to store neighborhood knowledge heavily relies on the availability of a sufficiently large dataset.
Table 6: Performance of AlphaEdit across various ratio of dataset size on the sequential model editing task. Eff., Gen. and Spe. denote Efficacy, Generalization and Specificity, respectively.
Model Ratio Counterfact ZsRE
Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$
LLaMA3 1.0 98.90$\pm$1.21 94.22$\pm$0.89 67.88$\pm$1.34 94.47$\pm$0.97 91.13$\pm$1.02 32.55$\pm$1.78
0.9 98.32$\pm$0.92 93.87$\pm$1.56 66.23$\pm$1.18 94.12$\pm$1.44 91.76$\pm$1.02 31.89$\pm$1.23
0.8 96.75$\pm$1.35 92.45$\pm$0.78 66.45$\pm$0.99 94.12$\pm$1.23 90.95$\pm$1.42 30.67$\pm$1.09
0.7 95.66$\pm$0.76 93.11$\pm$0.98 64.89$\pm$1.41 93.87$\pm$1.11 90.45$\pm$0.97 29.34$\pm$0.84
0.6 96.12$\pm$0.86 91.34$\pm$1.23 63.34$\pm$0.94 93.87$\pm$1.36 91.12$\pm$1.11 29.34$\pm$1.25
0.5 97.93$\pm$1.23 94.01$\pm$1.09 63.51$\pm$0.97 92.96$\pm$0.89 91.67$\pm$1.03 28.56$\pm$1.34
0.4 95.88$\pm$0.78 92.67$\pm$1.11 61.78$\pm$1.09 92.98$\pm$1.09 91.76$\pm$1.28 28.56$\pm$0.99
0.3 96.98$\pm$1.67 93.22$\pm$0.99 58.56$\pm$1.08 93.12$\pm$1.43 89.12$\pm$1.23 27.56$\pm$1.67
0.2 97.45$\pm$0.97 91.89$\pm$1.22 56.89$\pm$0.89 93.01$\pm$0.84 89.99$\pm$1.09 26.34$\pm$1.34
0.1 95.21$\pm$1.03 90.12$\pm$1.45 56.12$\pm$1.22 92.01$\pm$1.02 89.97$\pm$1.28 25.89$\pm$1.47
GPT2-XL 1.0 99.50$\pm$0.98 93.95$\pm$1.13 66.39$\pm$0.89 94.81$\pm$1.56 86.11$\pm$1.24 25.88$\pm$1.42
0.9 97.82$\pm$1.43 92.78$\pm$0.87 65.24$\pm$0.92 93.67$\pm$1.05 85.73$\pm$1.12 25.05$\pm$1.32
0.8 98.47$\pm$1.12 92.54$\pm$1.32 64.89$\pm$0.76 93.21$\pm$0.99 85.48$\pm$0.78 23.98$\pm$1.24
0.7 96.23$\pm$1.09 93.21$\pm$1.24 64.45$\pm$0.98 94.05$\pm$1.09 85.74$\pm$0.89 23.67$\pm$1.11
0.6 99.12$\pm$0.87 91.33$\pm$1.07 64.45$\pm$0.93 94.31$\pm$0.99 84.85$\pm$1.45 23.45$\pm$0.98
0.5 95.68$\pm$0.92 90.89$\pm$1.34 61.76$\pm$0.76 94.74$\pm$0.87 85.92$\pm$1.23 22.78$\pm$0.76
0.4 97.54$\pm$1.45 91.76$\pm$1.23 60.23$\pm$1.09 93.52$\pm$0.98 84.45$\pm$0.88 22.34$\pm$1.01
0.3 99.01$\pm$1.09 90.32$\pm$1.11 58.92$\pm$0.92 93.01$\pm$1.34 83.78$\pm$0.99 22.67$\pm$1.21
0.2 95.89$\pm$0.78 91.03$\pm$1.03 58.14$\pm$1.23 93.04$\pm$1.09 85.07$\pm$0.95 22.45$\pm$1.15
0.1 96.35$\pm$1.02 91.03$\pm$1.32 58.14$\pm$1.14 93.04$\pm$1.22 85.07$\pm$1.43 21.11$\pm$1.12
GPT-J 1.0 99.75$\pm$1.15 96.38$\pm$1.45 75.48$\pm$1.25 99.79$\pm$1.28 96.00$\pm$1.67 28.29$\pm$1.32
0.9 99.43$\pm$1.09 96.14$\pm$1.08 74.75$\pm$1.11 97.63$\pm$1.03 96.11$\pm$0.98 27.12$\pm$1.43
0.8 98.34$\pm$0.76 96.03$\pm$1.12 75.21$\pm$0.89 97.65$\pm$0.87 96.01$\pm$1.32 25.89$\pm$0.89
0.7 98.21$\pm$1.23 95.11$\pm$1.56 73.58$\pm$1.01 98.78$\pm$1.15 95.34$\pm$0.94 25.09$\pm$1.09
0.6 98.94$\pm$1.17 95.33$\pm$1.12 73.89$\pm$1.21 99.05$\pm$1.09 95.45$\pm$0.84 24.87$\pm$1.45
0.5 98.46$\pm$1.21 94.89$\pm$1.09 70.88$\pm$1.05 98.94$\pm$1.04 95.12$\pm$1.28 23.78$\pm$0.97
0.4 97.74$\pm$0.88 94.76$\pm$1.04 69.89$\pm$1.18 98.31$\pm$1.23 94.91$\pm$1.09 23.67$\pm$1.22
0.3 96.74$\pm$1.32 94.34$\pm$0.95 67.95$\pm$0.84 97.58$\pm$0.98 94.23$\pm$1.15 22.78$\pm$1.12
0.2 96.94$\pm$1.09 94.73$\pm$1.24 68.04$\pm$0.92 97.34$\pm$0.97 94.12$\pm$1.02 23.45$\pm$1.09
0.1 97.02$\pm$0.89 94.73$\pm$0.98 68.04$\pm$1.09 97.58$\pm$1.21 94.23$\pm$0.89 23.67$\pm$1.32
Figure 11: Performance of AlphaEdit across various ratio of dataset size on the sequential model editing task. Best viewed in color.
Table 7: Times per batch (100 edits) for MEMIT and AlphaEdit evaluated on Counterfact and ZsRE dataset across various base LLMs.
Method Counterfact ZsRE
LLaMA3 GPT-J GPT2-XL LLaMA3 GPT-J GPT2-XL
MEMIT 222.51s 334.74s 474.14s 231.32s 344.21s 488.37s
AlphaEdit 223.24s 334.93s 476.79s 231.40s 345.52s 490.25s
C.9 Runtime Evaluation of AlphaEdit
The computational complexity of the null-space projection in AlphaEdit depends solely on the dimension of the hidden dimension $d_{0}$ within the base LLM. Specifically, calculating the null-space projection matrix only requires operations on ${\bm{K}}_{0}{\bm{K}}_{0}^{T}\in\mathbb{R}^{d_{0}\times d_{0}}$, which are independent of the number of layers, model size, or the knowledge base size.
To empirically validate the scalability of our method, we measured the average runtime for performing 100 edits with AlphaEdit and MEMIT on three LLMs with different model sizes and knowledge bases: LLaMA3, GPT-J, and GPT2-XL. The results are summarized in Table 7.
From these results, it is evident that AlphaEdit does not incur additional runtime overhead compared to MEMIT, even as the model size or knowledge base grows. This supports our claim that the null-space projection method is highly scalable and practical for large-scale model editing tasks.
Appendix D Visualizing the Counterfact and ZSRE Datasets Through Examples